# Out-of-time ordered correlation functions for the localized $f$ electrons in the Falicov-Kimball model

### v1

@@ VII Conclusions
We have examined how an indirect retarded interaction between qubits can affect the scrambling of quantum information. We considered the dynamics of the localized two-level states ($f$-electrons) interacting with the itinerant $d$-electrons in the Falicov-Kimball model and found exact results for the OTOCs constructed via state-flip operators ($f$-electron creation and annihilation operators) within dynamical mean field theory. Despite the classical (Ising) nature of our two-level states (the occupation $n_{f}$ is a conserved quantity), their dynamics is nontrivial and is governed by the winding of the characteristic functions for the corresponding Toeplitz matrices. With an increase of the local interaction potential $U$ between $f$ and $d$ electrons, the Lyapunov exponent (logarithmic derivative) of the OTOCs start to display oscillations at high temperatures similar to those seen in quantum spin models. On the other hand, at low temperatures Fermi statistics and screening effects related to the orthogonality catastrophe become important, which strongly affects the OTOC’s behavior. With a further increase of the interaction $U$, we find a fast damping of these oscillations due to the effects of disorder and, on the other hand, makes the OTOC’s “density-of-states” (moments of the Kirkwood-Dirac quasiprobabilities) become negative, which is another manifestation of the quantum information scrambling.
@@ Appendix A Winding of characteristic function for one-particle propagator
$\lim_{N\rightarrow\infty}D_{N}=\exp\left[Ng_{0}+\sum_{n=1}^{\infty}ng_{n}g_{-n}\right]$ (77)
@@
$\lim_{t\rightarrow\infty}D(t)=\exp\left[\frac{t}{2\pi}\int_{-\infty}^{\infty}\!\!\mathrm{d}\omega\ln C(\omega)+\int_{0}^{\infty}\!\!\mathrm{d}t^{\prime}t^{\prime}g(t^{\prime})g(-t^{\prime})\right],$ (81)
@@ Appendix B Winding of characteristic function for OTOC
Appendix C Winding of the Lyapunov exponent
Figure A.6: Winding of the Lyapunov exponent $L_{0}(t)$ for (a) $U=0.4$, (b) $U=0.9$, (c) $U=2.0$, $n_{f}=0.5$, and $n_{d}=0.5$.
The existence of three different winding numbers for the characteristic function is easily seen on the Nyquist plots for the Lyapunov exponent (73) too, see Fig. A.6 for an example. For $U<0.44$, the Lyapunov exponent smoothly evolves to some steady value. For $0.44<U<0.866$ and at high temperatures, $L_{0}(t)$ first displays evolution similar to what was seen in the previous case. It then is replaced by circulations, whereas at low temperatures we still observe a smooth evolution to a steady value. For $0.866<U<1.235$, we observe circulations of the Lyapunov exponent at high temperatures, which are destroyed (become more chaotic) with decreasing temperature and are replaced by some complicated looped trajectories. For $U>1.235$, due to the presence of pseudogap and Mott gap, the trajectories become even more chaotic, Fig. A.6.

### v2

@@ VII Conclusions
We have examined how an indirect retarded interaction between qubits can affect the scrambling of quantum information. We considered the dynamics of the localized two-level states ($f$-electrons) interacting with the itinerant $d$-electrons in the Falicov-Kimball model and found exact results for the OTOCs constructed via state-flip operators ($f$-electron creation and annihilation operators) within dynamical mean field theory. Despite the classical (Ising) nature of our two-level states (the occupation $n_{f}$ is a conserved quantity), their dynamics is nontrivial and is governed by the winding of the characteristic functions for the corresponding Toeplitz matrices. With an increase of the local interaction potential $U$ between $f$ and $d$ electrons, the logarithmic derivative of the OTOCs start to display oscillations at high temperatures similar to those seen in quantum spin models. On the other hand, at low temperatures Fermi statistics and screening effects related to the orthogonality catastrophe become important, which strongly affects the OTOC’s behavior. With a further increase of the interaction $U$, we find a fast damping of these oscillations due to the effects of disorder and, on the other hand, makes the OTOC’s “density-of-states” (moments of the Kirkwood-Dirac quasiprobabilities) become negative, which is another manifestation of the quantum information scrambling.
Our results were obtained at half filling with $n_{d}=n_{f}=0.5$, but our main conclusions remain valid in when $n_{f}+n_{d}=1$ holds, with the remark that now asymptotic behavior of each contribution $F_{0}(t)$ and $F_{1}(t)$ in (57) will be different with different windings of the corresponding Toeplitz matrices. On the other hand, in the doped case, with $n_{f}+n_{d}\neq 1$, our system will be in a metallic phase and the low temperature behavior will be dominated by the localization singularities given by the Anderson orthogonality catastrophe, whereas at high temperatures the results will be similar to the one for un-doped case.
@@ Acknowledgements.
Data Availability
The data that support the findings of this article are openly available [46].
@@ Appendix A Winding of characteristic function for one-particle propagator
$\lim_{N\rightarrow\infty}D_{N}=\exp\Bigl[Ng_{0}+\sum_{n=1}^{\infty}ng_{n}g_{-n}\Bigr]$ (79)
@@
$\lim_{t\rightarrow\infty}D(t)=\exp\Bigl[\frac{t}{2\pi}\int_{-\infty}^{\infty}\!\!\mathrm{d}\omega\ln C(\omega)+\int_{0}^{\infty}\!\!\mathrm{d}t^{\prime}t^{\prime}g(t^{\prime})g(-t^{\prime})\Bigr],$ (83)
@@ Appendix B Winding of characteristic function for OTOC
Appendix C Winding of the logarithmic derivative
Figure A.6: Winding of the logarithmic derivative $L_{0}(t)$ for (a) $U=0.4$, (b) $U=0.9$, (c) $U=2.0$, $n_{f}=0.5$, and $n_{d}=0.5$.
The existence of three different winding numbers for the characteristic function is easily seen on the Nyquist plots for the logarithmic derivative (73) too, see Fig. A.6 for an example. For $U<0.44$, the logarithmic derivative smoothly evolves to some steady value. For $0.44<U<0.866$ and at high temperatures, $L_{0}(t)$ first displays evolution similar to what was seen in the previous case. It then is replaced by circulations, whereas at low temperatures we still observe a smooth evolution to a steady value. For $0.866<U<1.235$, we observe circulations of the logarithmic derivative at high temperatures, which are destroyed (become more chaotic) with decreasing temperature and are replaced by some complicated looped trajectories. For $U>1.235$, due to the presence of pseudogap and Mott gap, the trajectories become even more chaotic, Fig. A.6.